# Hermite trace polynomials and chaos decompositions for the Hermitian Brownian motion

### v2

@@ 6.2. The isomorphisms
For a partition $\mu=1^{k_{1}}2^{k_{2}}\ldots m^{k_{m}}$, a (long) computation (see [AGH+17] for a similar computation in the unitary case) shows that
$D^{\ast}(p_{\mu})=p_{\mu}\Biggl{[}2k_{2}\frac{N^{2}}{p_{2}}+2N\sum_{l=3}^{m}k_{l}l\frac{p_{l-2}}{p_{l}}+\sum_{l=2}^{m}k_{l}(k_{l}-1)l^{2}\frac{p_{2l-2}}{p_{l}^{2}}+k_{1}(k_{1}-1)\frac{N}{p_{1}^{2}}\\ +\sum_{2\leq l_{1}<l_{2}\leq m}2k_{l_{1}}k_{l_{2}}l_{1}l_{2}\frac{p_{l_{1}+l_{2}-2}}{p_{l_{1}}p_{l_{2}}}+2\sum_{l=2}^{m}k_{1}k_{l}l\frac{p_{l-1}}{p_{1}p_{l}}+\sum_{l=3}^{m}k_{l}l\sum_{a=1}^{l-3}\frac{p_{a}p_{l-2-a}}{p_{l}}\Biggr{]}$
On the other hand, for a permutation $\alpha$ of cycle type $\mu=1^{k_{1}}2^{k_{2}}\ldots m^{k_{m}}$, a contraction by a transposition
  • Eliminates a cycle of length $2$: in $k_{2}$ cases, with weight $q^{-1}$.
  • Turns a cycle of length $l>2$ into a cycle of length $l-2$: in $k_{l}l$ cases, with weight $1$.
  • Splits a cycle of length $l$ into those of different lengths $a<l-2-a$: in $k_{l}l$ cases, with weight $q$.
  • Splits a cycle of length $l$ into those of equal lengths $(l-2)/2$: in $k_{l}l/2$ cases, with weight $q$.
@@
  • Eliminates a cycle of length $1$ and turns a cycle of length $l$ into a cycle of length $l-1$: in $k_{1}k_{2}l$ cases, with weight $q$.
  • Glues together two cycles of different lengths $2\leq l_{1}<l_{2}$: in $k_{l_{1}}l_{l_{2}}l_{1}l_{2}$ cases, with weight $q$.
  • Glues together two cycles of equal length $l\geq 2$: in $k_{l}(k_{l}-1)l^{2}/2$ cases, with weight $q$.
Comparing these two results, we see that $D^{\ast}p_{\mu}$ corresponds to $2N\mathcal{E}[\mathnormal{T}\left(\mathcal{L}\mu\right)]$, giving the first relation. The second relation follows from $E^{\ast}p_{\mu}=np_{\mu}$. ∎
@@
Since their leading terms differ by a factor of $\frac{\left|\lambda\right|!}{c_{\lambda}n!}$, it suffices to verify that $\mathcal{E}[\mathnormal{W}\left(\chi^{\lambda}\otimes h^{\otimes n}\right)]$ is an eigenfunction of $D^{\ast}-E^{\ast}$ with eigenvalue $-n$. Indeed, for $\left\|h\right\|^{2}=\frac{1}{N}$,
@@ 6.2.2. $q=0$
We will now denote $\alpha\otimes_{s}F$ by $\tilde{W}(\alpha\otimes_{s}F)$. For $\ell=\left|\pi\right|$, define the normalized contractions
@@
$\tilde{W}(\alpha\otimes_{s}F)=\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}(-1)^{n-\left|\pi\right|}\tilde{T}(\tilde{C}_{\pi}(\alpha)\otimes_{s}C_{\pi}(F))$
@@
Finally, define
@@
If we assume that $\tilde{T}(\alpha\otimes F)=q^{\operatorname{\mathrm{cyc}}_{0}(\alpha)}\mathnormal{T}\left(\alpha\otimes F\right)$, then for $q=0$, such a multiple is zero unless $\alpha$ is a single cycle (containing 00). As seen below, this need not be the case in general.
@@
Let $\lambda$ be an interval partition of $[0,n]$. As discussed in Section 2.2, each permutation $\alpha\in S_{0}(n)$ is conjugate, under the action of $S(n)$, to a permutation with cycle structure $\lambda$ in which the elements in each cycle, as well as the cycles, appear in increasing order. Such permutation is not unique. In the results below, we will only consider permutations of this type; the results are easily extended to general $\alpha\in S_{0}(n)$, but the notation gets heavier.
@@
Let $\mathcal{F}_{f}(\mathcal{H})=\mathbb{C}\Omega\oplus\bigoplus_{n=1}^{\infty}\overline{\mathcal{H}^{\otimes n}}$ be the full Fock space of $\mathcal{H}_{\mathbb{R}}$. For $f\in\mathcal{H}_{\mathbb{R}}$, denote by $S(f)$ the semicircular element corresponding to $f$ in its standard representation on $\mathcal{F}_{f}(\mathcal{H}_{\mathbb{R}})$. Denote by $\Gamma_{f}(\mathcal{H}_{\mathbb{R}})$ the algebra generated by $\left\{X(h):h\in\mathcal{H}_{\mathbb{R}}\right\}$, and by $\Phi$ the vacuum state on it. Denote by $U(h_{1}\otimes\ldots\otimes h_{n})$ the multivariate Chebyshev polynomials, which are the elements of $\Gamma_{f}(\mathcal{H}_{\mathbb{R}})$ which satisfy $U(h_{1}\otimes\ldots\otimes h_{n})\Omega=h_{1}\otimes\ldots\otimes h_{n}$. They are also determined by the recursion
$S(f)U(h_{1}\otimes\ldots\otimes h_{n})=U(f\otimes h_{1}\otimes\ldots\otimes h_{n})+\left\langle f,h_{1}\right\rangle U(h_{2}\otimes\ldots\otimes h_{n}).$
@@
Let $q=0$, and $\lambda,\mu\in S_{0}(n)$ as in Remark 6.21.
@@
    $\mathcal{E}_{0}[\tilde{T}(\alpha\otimes(h_{1}\otimes\ldots\otimes h_{n}))]=\Phi_{\alpha}[S(h_{1}),\ldots,S(h_{n})].$
    Then $\mathcal{E}_{0}$ extends to a homomorphism from $\overline{TP}(\mathcal{H}_{\mathbb{R}})$, and to an isometric isomorphism from $\mathcal{F}_{0}(\mathcal{H})$ to $L^{2}(\Gamma_{f}(\mathcal{H}_{\mathbb{R}}),\Phi)\simeq\mathcal{F}_{f}(\mathcal{H})$.
@@
    If $\lambda$ is not a single cycle (containing 00), then $\tilde{W}({\lambda}\otimes F)=0$. Otherwise,
@@
Note first that
@@
Moreover, for $\alpha=\lambda$ as in Remark 6.21, this is equal to 00 if and only if $\pi\leq\lambda$, and on each block of $\lambda$, $\pi$ is non-crossing. Therefore
$\tilde{T}({\lambda}\otimes F)=\sum_{\begin{subarray}{c}\pi\in\mathcal{NC}_{1,2}(n)\\ \pi\leq{\lambda}\end{subarray}}\tilde{W}(\tilde{C}_{\pi}({\lambda})\otimes C_{\pi}(F)),$
@@
$\varphi\left[\tilde{T}({\lambda}\otimes F)\right]=\sum_{\begin{subarray}{c}\pi\in\mathcal{NC}_{2}(n)\\ \pi\leq{\lambda}\end{subarray}}C_{\pi}(F)=\Phi[\Phi_{\lambda}[S(h_{1}),\ldots,S(h_{n})]].$
The homomorphism property follows as before, so $\mathcal{E}_{0}$ is an isometric isomorphism. In particular, if $(0)$ is a cycle in $\lambda$, then
(13) $\tilde{T}({\lambda}\otimes F)=\varphi\left[\tilde{T}({\lambda}\otimes F)\right]=\sum_{\begin{subarray}{c}\pi\in\mathcal{NC}_{2}(n)\\ \pi\leq{\lambda}\end{subarray}}C_{\pi}(F)$
is a scalar. Next,
$\tilde{W}(\lambda\otimes F)=\sum_{\begin{subarray}{c}\pi\in\mathcal{NC}_{1,2}(n)\\ \pi\leq\lambda\end{subarray}}(-1)^{n-\left|\pi\right|}\tilde{T}(\tilde{C}_{\pi}({\lambda})\otimes C_{\pi}(F)).$
Suppose $\lambda$ contains a cycle $V$ not containing 00. Then the sum above is
$\begin{split}\tilde{W}(\alpha\otimes F)&=\sum_{\begin{subarray}{c}\pi\in\mathcal{NC}_{1,2}([n]\setminus V)\\ \pi\leq\lambda\end{subarray}}(-1)^{n-\left|V\right|-\left|\pi\right|}\tilde{T}(\tilde{C}_{\pi}({\lambda|_{V^{c}}})\otimes C_{\pi}(\bigotimes_{i\in[n]\setminus V}h_{i}))\\ &\qquad\cdot\sum_{\tau\in\mathcal{NC}_{1,2}(V)}(-1)^{\left|V\right|-\left|\tau\right|}\tilde{T}(\tilde{C}_{\tau}({V})\otimes C_{\tau}(\bigotimes_{i\in V}h_{i})).\end{split}$
Since $V$ does not contain 00, denoting $F_{V}=\bigotimes_{i\in V}h_{i}$ and using (13), the second of these sums is
$\sum_{\tau\in\mathcal{NC}_{1,2}(V)}(-1)^{\left|V\right|-\left|\tau\right|}\sum_{\sigma\in\mathcal{NC}_{2}(\mathrm{Sing}\left(\tau\right))}C_{\sigma}C_{\tau}F_{V}=\sum_{\rho\in\mathcal{NC}_{2}(V)}\sum_{S\in\mathrm{Pair}\left(\rho\right)}(-1)^{\left|V\right|-\left|S\right|}C_{\rho}F_{V}=0.$
On the other hand, $\mathcal{E}_{0}[\tilde{W}({(01\ldots n)}\otimes(h_{1}\otimes\ldots\otimes h_{n}))]$ is a polynomial in $\left\{S(h_{1}),\ldots,S(h_{n})\right\}$, with leading term $S(h_{1})\ldots S(h_{n})$, and
$\begin{split}&\Phi[\mathcal{E}_{0}[\tilde{W}({(01\ldots k)}\otimes(g_{1}\otimes\ldots\otimes g_{k}))]^{\ast}\mathcal{E}_{0}[\tilde{W}({(01\ldots n)}\otimes(h_{1}\otimes\ldots\otimes h_{n}))]]\\ &\quad=\varphi[\tilde{W}({(01\ldots k)}\otimes(g_{1}\otimes\ldots\otimes g_{k}))^{\ast}\tilde{W}({(01\ldots n)}\otimes(h_{1}\otimes\ldots\otimes h_{n}))]\\ &\quad=\delta_{n=k}\prod_{i=1}^{n}\left\langle h_{i},g_{i}\right\rangle\\ &\quad=\Phi[U(g_{1}\otimes\ldots\otimes g_{k})^{\ast}U(h_{1}\otimes\ldots\otimes h_{n})]\end{split}$
since there is only one non-crossing pair partition in $\mathcal{P}(n,n)$. Since $\mathcal{E}_{0}$ is an isometry, the final statement follows. ∎
Funding
This work was supported in part by a Simons Foundation Collaboration Grant.

### v3

@@ 6.2. The isomorphisms
By linearity, it suffices to prove the result for $\eta=\alpha\in S(n)$ a single permutation. Suppose the cycle structure of $\alpha$ is given by the partition $\mu=1^{k_{1}}2^{k_{2}}\ldots m^{k_{m}}$. The second relation follows from $E^{\ast}p_{\mu}=np_{\mu}$. Recall also that for a permutation $\alpha$, the Laplacian $\mathcal{L}(\alpha)$ is the sum of contractions of $\alpha$ with respect to all transpositions. Thus the first relation reduces to showing that
$D^{\ast}p_{\mu}(x_{1},\ldots,x_{N})=2N\sum_{\begin{subarray}{c}\tau\in S(n)\\ \text{ a transposition}\end{subarray}}p_{\widetilde{C_{\tau}(\alpha)}}(x_{1},\ldots,x_{N}),$
where $\widetilde{C_{\tau}(\alpha)}$ is the cycle type partition of $C_{\tau}(\alpha)$. For the left-hand side, a (long) computation (see [1] for a similar computation in the unitary case) shows that
$D^{\ast}(p_{\mu})=2p_{\mu}\Biggl[k_{2}\frac{N^{2}}{p_{2}}+N\sum_{l=3}^{m}k_{l}l\frac{p_{l-2}}{p_{l}}+\sum_{l=2}^{m}\frac{k_{l}(k_{l}-1)l^{2}}{2}\frac{p_{2l-2}}{p_{l}^{2}}+\frac{k_{1}(k_{1}-1)}{2}\frac{N}{p_{1}^{2}}\\ +\sum_{2\leq l_{1}<l_{2}\leq m}k_{l_{1}}k_{l_{2}}l_{1}l_{2}\frac{p_{l_{1}+l_{2}-2}}{p_{l_{1}}p_{l_{2}}}+\sum_{l=2}^{m}k_{1}k_{l}l\frac{p_{l-1}}{p_{1}p_{l}}+\sum_{l=3}^{m}\frac{k_{l}l}{2}\sum_{a=1}^{l-3}\frac{p_{a}p_{l-2-a}}{p_{l}}\Biggr]$
For the right-hand side, for a permutation $\alpha$ of cycle type $\mu=1^{k_{1}}2^{k_{2}}\ldots m^{k_{m}}$ and $q=1/N$, a contraction by a transposition
  • Eliminates a cycle of length $2$: in $k_{2}$ cases, with weight $N$.
  • Turns a cycle of length $l>2$ into a cycle of length l2l-2: in $k_{l}l$ cases, with weight $1$.
  • Glues together two cycles of equal length $l\geq 2$: in $k_{l}(k_{l}-1)l^{2}/2$ cases, with weight $N^{-1}$.
@@
  • Glues together two cycles of different lengths $2\leq l_{1}<l_{2}$: in $k_{l_{1}}l_{l_{2}}l_{1}l_{2}$ cases, with weight $N^{-1}$.
  • Eliminates a cycle of length $1$ and turns a cycle of length $l$ into a cycle of length l1l-1: in $k_{1}k_{2}l$ cases, with weight $N^{-1}$.
  • Splits a cycle of length $l$ into those of different lengths a<l2aa<l-2-a: in $k_{l}l$ cases, with weight $N^{-1}$.
  • Splits a cycle of length $l$ into those of equal lengths (l2)/2(l-2)/2: in $k_{l}l/2$ cases, with weight $N^{-1}$.
The expressions match up term-by-term. ∎
@@
Since their leading terms differ by a factor of $\frac{\left|\lambda\right|!}{c_{\lambda}n!}$, it suffices to verify that $\mathcal{E}[\mathnormal{W}\left(\chi^{\lambda}\otimes h^{\otimes n}\right)]$ is an eigenfunction of $D^{\ast}-E^{\ast}$ with eigenvalue n-n. Indeed, for $\left\|h\right\|^{2}=\frac{1}{N}$,
@@ 6.2.2. The case of $q=0$
In this section, we will denote $\alpha\otimes_{s}F$ by $\tilde{W}(\alpha\otimes_{s}F)$. For $\ell=\left|\mathrm{Pair}\left(\pi\right)\right|=\left|\pi\right|$, define the normalized contractions
@@
(20) $\tilde{W}(\alpha\otimes_{s}F)=\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}(-1)^{\left|\pi\right|}\tilde{T}(\tilde{C}_{\pi}(\alpha)\otimes_{s}C_{\pi}(F))$
@@
Finally, define the linear functional $\varphi$ by
@@
Then for $\alpha,\beta\in S_{0}(n)$, Theorem 5.15 gets updated to
(21) $\varphi\left[\tilde{W}(\beta\otimes_{s}G)^{\ast}\tilde{W}(\alpha\otimes_{s}F)\right]=q^{\operatorname{\mathrm{cyc}}_{0}(\alpha)+\operatorname{\mathrm{cyc}}_{0}(\beta)}\sum_{\sigma\in S(n)}q^{\left|\sigma^{-1}\alpha\sigma\beta^{-1}\right|}\left\langle F,U_{\sigma}G\right\rangle.$
Also for $\alpha\in S_{0}(2n)$, that theorem gives
$\varphi\left[\tilde{T}(\alpha\otimes F)\right]=\sum_{\pi\in\operatorname{\mathcal{P}}_{2}(2n)}q^{\left|\pi\alpha\right|-n+\operatorname{\mathrm{cyc}}_{0}(\alpha)}C_{\pi}(F)=\sum_{\pi\in\operatorname{\mathcal{P}}_{2}(2n)}q^{\left|\pi\alpha\right|+n-\left|\alpha\right|}C_{\pi}(F).$
@@
If we assume that $\tilde{T}(\alpha\otimes F)=q^{\operatorname{\mathrm{cyc}}_{0}(\alpha)}\mathnormal{T}\left(\alpha\otimes F\right)$, then for $q=0$, such a multiple is zero unless $\alpha$ is a single cycle (containing 0). The definition above is a little more flexible.
@@
Let $\lambda$ be an interval partition of $[0,n]$. As discussed in Section 2.2, each permutation $\alpha\in S_{0}(n)$ is conjugate, under the action of $S(n)$, to a permutation with cycle structure $\lambda$ in which the elements in each cycle, as well as the cycles, appear in increasing order. Such a permutation is not unique. In the results below, we will only consider permutations of this type; the results are easily extended to general $\alpha\in S_{0}(n)$, but the notation gets heavier.
@@
Let $\mathcal{F}_{f}(\mathcal{H})=\mathbb{C}\Omega\oplus\bigoplus_{n=1}^{\infty}\overline{\mathcal{H}^{\otimes n}}$ be the full Fock space of $\mathcal{H}_{\mathbb{R}}$. For $h\in\mathcal{H}_{\mathbb{R}}$, denote by $S(h)$ the semicircular element corresponding to $h$ in its standard representation on $\mathcal{F}_{f}(\mathcal{H}_{\mathbb{R}})$. Denote by $\Gamma_{f}(\mathcal{H}_{\mathbb{R}})$ the algebra generated by $\left\{X(h):h\in\mathcal{H}_{\mathbb{R}}\right\}$, and by $\Phi$ the vacuum state on it. Denote by $U(h_{1}\otimes\ldots\otimes h_{n})$ the multivariate Chebyshev polynomials (of the second kind), which are the elements of $\Gamma_{f}(\mathcal{H}_{\mathbb{R}})$ that satisfy $U(h_{1}\otimes\ldots\otimes h_{n})\Omega=h_{1}\otimes\ldots\otimes h_{n}$. They are also determined by the recursion
(22) $U(f\otimes h_{1}\otimes\ldots\otimes h_{n})=S(f)U(h_{1}\otimes\ldots\otimes h_{n})-\left\langle f,h_{1}\right\rangle U(h_{2}\otimes\ldots\otimes h_{n}).$
@@
See Lecture 7 of [32] and Section 5 of [5] for additional background on the full Fock space, Chebyshev polynomials, and free analysis.
@@
Let $q=0$, and $\lambda\in S_{0}(n)$ as in Remark 6.21. Denote $V_{0}$ the block of $\lambda$ that contains 0.
@@
    (23) $\mathcal{E}_{0}[\tilde{T}(\lambda\otimes(h_{1}\otimes\ldots\otimes h_{n}))]\\ =\Phi_{\lambda}[S(h_{1}),\ldots,S(h_{n})]=S(h_{1})\ldots S(h_{\left|V_{0}\right|})\prod_{\begin{subarray}{c}V\in\lambda\\ V\neq V_{0}\end{subarray}}\sum_{\pi\in\mathcal{NC}_{2}(V)}C_{\pi}(\bigotimes_{i\in V}h_{i}),$
    where the notation $\Phi_{\lambda}$ is defined at the beginning of Section 6. Then $\mathcal{E}_{0}$ extends to a homomorphism from $\overline{\mathcal{TP}}(\mathcal{H}_{\mathbb{R}})$, and to an isometric isomorphism from $\mathcal{F}_{0}(\mathcal{H})$ to $L^{2}(\Gamma_{f}(\mathcal{H}_{\mathbb{R}}),\Phi)\simeq\mathcal{F}_{f}(\mathcal{H})$.
@@
    If $\lambda$ is not a single cycle (containing 0), then $\mathcal{E}_{0}[\tilde{W}({\lambda}\otimes F)]=0$. Otherwise,
@@
Note first that for $\ell=\left|\pi\right|$,
@@
Moreover, for $\alpha=\lambda$ as in Remark 6.21, this is equal to 0 if and only if $\pi\leq\lambda$, and on each block of $\lambda$, $\pi$ is non-crossing (see Lecture 23 of [32]). Therefore equation (19) reduces to
(24) $\tilde{T}({\lambda}\otimes F)=\sum_{\begin{subarray}{c}\pi\in\mathcal{NC}_{1,2}(n)\\ \pi\leq{\lambda}\end{subarray}}\tilde{W}(\tilde{C}_{\pi}({\lambda})\otimes C_{\pi}(F)),$
@@
$\varphi\left[\tilde{T}({\lambda}\otimes F)\right]=\sum_{\begin{subarray}{c}\pi\in\mathcal{NC}_{2}(n)\\ \pi\leq{\lambda}\end{subarray}}C_{\pi}(F)=\Phi[\Phi_{\lambda}[S(h_{1}),\ldots,S(h_{n})]],$
where the second equality follows from the properties of semicircular elements. The homomorphism property follows as before, so $\mathcal{E}_{0}$ is an isometric isomorphism. In particular, to prove the second equality in (23), it suffices to note that if $(0)$ is a cycle in $\lambda$, then
(25) $\mathcal{E}_{0}[\tilde{T}({\lambda}\otimes F)]=\Phi_{\lambda}[S(h_{1}),\ldots,S(h_{n})]=\sum_{\begin{subarray}{c}\pi\in\mathcal{NC}_{2}(n)\\ \pi\leq{\lambda}\end{subarray}}C_{\pi}(F)$
is a scalar. Next, for $q=0$, the relation (21) and the isometry property imply that $\mathcal{E}_{0}[\tilde{W}({\lambda}\otimes F)]=0$ whenever $\operatorname{\mathrm{cyc}}_{0}(\lambda)>0$, that is, if $\lambda$ is not a single cycle containing 0. Finally, $\mathcal{E}_{0}[\tilde{W}({(01\ldots n)}\otimes(h_{1}\otimes\ldots\otimes h_{n}))]$ is a polynomial in $\left\{S(h_{1}),\ldots,S(h_{n})\right\}$, with the leading term $S(h_{1})\ldots S(h_{n})$, and
$\begin{split}&S(f)\mathcal{E}_{0}[\tilde{W}((01\ldots n)\otimes(h_{1}\otimes\ldots\otimes h_{n}))]\\ &\quad=\mathcal{E}_{0}[\tilde{W}((01\ldots n(n+1))\otimes(f\otimes h_{1}\otimes\ldots\otimes h_{n}))]\\ &\quad\quad-\sum_{k=1}^{n}\left\langle f,h_{k}\right\rangle\mathcal{E}_{0}[\tilde{W}((0k\ldots(n-1))(1\ldots(k-1))\otimes(f\otimes h_{1}\otimes\ldots\otimes\hat{h}_{k}\otimes\ldots\otimes h_{n}))]\\ &\quad=\mathcal{E}_{0}[\tilde{W}((01\ldots n(n+1))\otimes(f\otimes h_{1}\otimes\ldots\otimes h_{n}))]\\ &\quad\quad-\left\langle f,h_{1}\right\rangle\mathcal{E}_{0}[\tilde{W}((01\ldots(n-1))\otimes(f\otimes h_{2}\otimes\ldots\otimes h_{n}))]\end{split}$
since the remaining terms in the sum are zero. Comparing with the recursion (22), we conclude that these are indeed the Chebyshev polynomials. ∎
Acknowledgments. The first author has discussed various aspects of this project with a number of people. He is especially grateful to Todd Kemp, Andu Nica, JM Landsberg, Marek Bożejko, and Jurij Volčič. The authors also thank the referees for an incredibly careful reading of the manuscript and numerous comments. Their suggestions resulted in substantial improvements, including the Gaussian Hilbert spaces approach and Theorem 6.11, as well as a better (in particular, correct) proof of Theorem 5.29. Both authors were supported by the Simons Foundation Collaboration Grant for Mathematicians #527486 .